# Murphy: Multi-Turn GRPO for Self Correcting Code Generation

### v1

@@ Appendix C Reflexion
Appendix D Implementation Details
We implement our framework on top of TRL (von Werra et al., 2020), which provides efficient distributed training and a modular implementation of GRPO. We integrate TRL with vLLM for fast inference and large-scale rollout execution, enabling scalable multi-turn training in our experiments. Prompts used to train Murphy are listed in App. E. All experiments use publicly available datasets. The base models (Qwen3, OLMo) are available for research use under the Apache 2.0 license.
D.1 Model Size and Compute Budget
@@
D.2 Hyperparameters
We set the KL regularization factor $\beta=0.04$, learning rate to $10^{-6}$, and weight decay to $0.1$ for both GRPO and Murphy  variants. Unless stated otherwise, the number of stages in Murphy is set to 2. For GRPO and the first stage of Murphy, we use 8 rollouts per prompt, while the second stage uses up to 8 rollouts per prompt (resulting in a maximum of 64 rollouts). To ensure a fair computational comparison, we also train GRPO with 72 rollouts. Following Yang et al., we set the temperature to 0.6 and top-$p$ to 0.95 for all experiments.
D.3 Package Parameters
@@
D.4 Additional Experiments
D.4.1 Ablation: Comparison with a Naive Multi-Turn Extension
D.4.2 Ablation: Effect of Training Dataset Size
D.4.3 Ablation: Effect of Pruning in Multi-Turn Murphy Training
@@
Table 4: Performance of Qwen3-1.7B and OLMo-2-1124-7B-Instruct variants on HumanEval, MBPP, and BigcodeBench benchmarks, reported as pass@1 accuracy (% mean $\pm$ stdev over 3 independent runs). Best results are bold, second-best are underlinedMurphy- MaRS outperforms Murphy  (Simple) on average.
@@
Murphy- MaRS, InterP $78.66\pm 2.66$ $84.15\pm 1.83$ $45.53\pm 1.17$ $61.20\pm 0.53$ $2.70\pm 0.68$ $7.43\pm 1.79$
Murphy- MaRS $79.27\pm 2.11$ $83.94\pm 1.41$ $44.80\pm 1.11$ $62.00\pm 1.40$ $1.35\pm 0.68$ $5.86\pm 0.39$
Appendix E Prompt Examples
@@
Appendix F Potential Risks
@@
Appendix G Icon Attributions

### v2

@@ Appendix E Reflexion
Appendix F Computational Cost of Multi-Turn vs. Single-Turn GRPO
We compare the computational cost of multi-turn Murphy training against single-turn GRPO under matched rollout and gradient budgets (72). Table 5 reports end-to-end FLOPs per GPU and per-step training latency computed using DeepSpeed’s (Rasley et al., 2020) flops-profiler [^4] https://www.deepspeed.ai/tutorials/flops-profiler/ for Qwen3-1.7B and Qwen3-4B. These measurements capture the full optimization step, including forward and backward passes, optimizer updates, and communication overhead.
In some cases, multi-turn Murphy may require fewer FLOPs per step than single-turn GRPO despite incorporating feedback into later-turn prompts. This reduction arises from the tree structure of multi-turn rollouts: only generations that fail to solve the problem are expanded in subsequent turns, while successful generations terminate early. Under a matched rollout budget, this early termination reduces the total number of tokens generated, outweighing the increased sequence length from feedback concatenation.
However, multi-turn training incurs moderately higher step latency (e.g., 1.34s vs. 1.15s for Qwen3-1.7B). This is because later-turn sequences are longer due to feedback concatenation, leading to increased memory bandwidth requirements and quadratically scaling attention costs even though fewer sequences are processed overall. The per-device batch size is set to 8 for Qwen3-1.7B and 4 for Qwen3-4B.
Table 5: End-to-end computational cost per optimization step for multi-turn Murphy and single-turn GRPO under matched rollout and gradient budgets.
Model Setting FLOPs / GPU Step Latency
Qwen3-1.7B Multi-turn (Murphy) 23.39 TFLOPs 1.34 s
Single-turn (GRPO) 27.69 TFLOPs 1.15 s
Qwen3-4B Multi-turn (Murphy) 10.49 TFLOPs 2.34 s
Single-turn (GRPO) 11.68 TFLOPs 2.20 s
Appendix G Implementation Details
We implement our framework on top of TRL (von Werra et al., 2020), which provides efficient distributed training and a modular implementation of GRPO. We integrate TRL with vLLM for fast inference and large-scale rollout execution, enabling scalable multi-turn training in our experiments. Prompts used to train Murphy are listed in App. H. All experiments use publicly available datasets. The base models (Qwen3, OLMo) are available for research use under the Apache 2.0 license.
G.1 Model Size and Compute Budget
@@
G.2 Hyperparameters
We set the KL regularization factor $\beta=0.04$, learning rate to $10^{-6}$, and weight decay to $0.1$ for both GRPO and Murphy  variants. Unless stated otherwise, the number of turns in Murphy is set to 2. For GRPO and the first turn of Murphy, we use 8 rollouts per prompt, while the second turn uses up to 8 rollouts per prompt (resulting in a maximum of 64 rollouts in turn 2). To ensure a fair computational comparison, we also train GRPO with 72 rollouts. Following Yang et al., we set the temperature to 0.6 and top-$p$ to 0.95 during inference for all experiments.
G.3 Package Parameters
@@
G.4 Additional Experiments
Table 6: Performance of Qwen3-1.7B and OLMo-2-1124-7B-Instruct variants on HumanEval, MBPP, and BigcodeBench benchmarks, reported as pass@1 accuracy (% mean $\pm$ stdev over 3 independent runs). Best results are bold, second-best are underlinedMurphy- MaRS outperforms Murphy (Simple) on average. Murphy (Simple) is “simple” in that it uses no fan-out: it performs 144 total rollouts by sampling 72 outputs in the first turn and then sampling another 72 outputs in a second turn that incorporates feedback from the first turn. We match the total number of rollouts/updates to ensure a fair comparison.
@@ G.4 Additional Experiments
G.4.1 Ablation: Comparison with a Naive Multi-Turn Extension
@@
G.4.2 Ablation: Effect of Training Dataset Size
G.4.3 Ablation: Effect of Pruning in Multi-Turn Murphy Training
@@
Murphy- MaRS, InterP $78.66\pm 2.66$ $\mathbf{84.15\pm 1.83}$ $\mathbf{45.53\pm 1.17}$ $61.20\pm 0.53$ $\mathbf{2.70\pm 0.68}$ $\mathbf{7.43\pm 1.79}$
Murphy- MaRS $\mathbf{79.27\pm 2.11}$ $83.94\pm 1.41$ $44.80\pm 1.11$ $\mathbf{62.00\pm 1.40}$ $1.35\pm 0.68$ $5.86\pm 0.39$
Appendix H Prompt Examples
@@
Appendix I Potential Risks
@@
Appendix J Icon Attributions